OpenStudy (anonymous):

The width of a rectangle is 7 inches less than 4 times the length. If x represents the length, write an algebraic expression in terms of x that represents the area of the rectangle. Simplify the expression

OpenStudy (anonymous):

\[4x^2-7x\]

OpenStudy (anonymous):

if x=length & y=width, y+7=4x right?

OpenStudy (anonymous):

A=xy=x(4x-7)

OpenStudy (anonymous):

it's rectangle

OpenStudy (angela210793):

w=4x-7 S=w*x=(4x-7)*x

OpenStudy (anonymous):

anand.stylish.007 - good answer :)

OpenStudy (anonymous):

thnq......